A discount of 17% is offered on an item previously costing $2.00. What is the discounted price?
Found 2 solutions by algebrahouse.com, stanbon:
Answer by algebrahouse.com(1659)

$2.00
x 0.17 {changed 17% to decimal by moving decimal point left two places}
-------
$0.34 discount {multiplied}

$2.00
- 0.34 {subtract discount from original price}
------
= $1.66 discounted price

For more help from me, visit: www.algebrahouse.com
Also, you can Follow @AlgebraHouse on Twitter


Answer by stanbon(75887)

A discount of 17% is offered on an item previously costing $2.00. What is the discounted price?
-----------
New cost:: 2.00 - 0.17*2.00 = 0.87*2.00 = $1.74
=============================
Cheers,
Stan H.